# A Family of Berndt-Type Integrals and Associated Barnes Multiple Zeta Functions

### v1

@@
Abstract. In this paper, we focus on calculating a specific class of Berndt integrals, which exclusively involves (hyperbolic) cosine functions. Initially, this integral is transformed into a Ramanujan-type hyperbolic (infinite) sum via contour integration. Subsequently, a function incorporating $\theta$ is defined. By employing the residue theorem, the mixed Ramanujan-type hyperbolic (infinite) sum with both hyperbolic cosine and hyperbolic sine in the denominator is converted into a simpler Ramanujan-type hyperbolic (infinite) sum, which contains only hyperbolic cosine or hyperbolic sine in the denominator. The simpler Ramanujan-type hyperbolic (infinite) sum is then evaluated using Jacobi elliptic functions, Fourier series expansions, and Maclaurin series expansions. Ultimately, the result is expressed as a rational polynomial of $\Gamma(1/4)$ and $\pi^{-1/2}$. Additionally, the integral is related to the Barnes multiple zeta function, which provides an alternative method for its calculation.
@@ 1 Introduction
where $s\geq 1$ and $m\geq 1$ for the “$+$’ case and $s\geq 2m+1$ otherwise. In fact, the study of Berndt-type integrals can be traced back to 1916, when the renowned Indian mathematician Ramanujan investigated the following improper integrals involving trigonometric and hyperbolic functions. He conjectured the integral identity (see [14, pp. 325-326])
@@
Here $K\equiv K(x)$ denotes the complete elliptic integral of the first kind and $K^{\prime}=K(1-x)$. Kuznetsov [12] generalized Ismail-type integrals using contour integration and theta functions, proving
$\displaystyle\frac{1}{2}\displaystyle\int_{-\infty}^{\infty}\displaystyle\frac{t^{n}\mathrm{d}t}{\cos(K\sqrt{t})+\cosh(K^{\prime}\sqrt{t})}=(-1)^{n}\displaystyle\frac{d^{2n+1}}{du^{2n+1}}\displaystyle\frac{{\rm sn}(u,k)}{{\rm cd}(u,k)}|_{u=0},$ (1.3)
@@
In the past decade, research on these analogous integrals has gradually attracted increasing interest, yielding notable progress. Recent advances in this direction can be found in the works of [7, 13, 16, 17, 21, 22, 24] and the references therein. A significant turning point in this line of research occurred in 2016 when Professor Berndt [6] developed a systematic approach to study ${\rm BI}^{\pm}(s,1)$. Berndt further advanced the field by evaluating such integrals directly using Cauchy’s residue theory, Fourier series expansions, and Maclaurin series expansions of Jacobi elliptic functions. In their series of works [17, 21, 22], Rui, Xu, and Zhao developed an extended framework based on Berndt’s approach, which led to fundamental structural theorems for higher-order Berndt-type integrals. Letting $X=\Gamma^{4}(1/4)$ and $Y=\pi^{-1}$, they proved that for all integers $m\geq 1$ and $p\geq[m/2]$ (see [22, Thms 1.1 and 1.2]),
@@
Very recently, a groundbreaking connection between Berndt-type integrals and Barnes multiple zeta functions (See Section 7) was established by Bradshaw and Vignat [7, Cor. 5], who also linked Jacobi elliptic functions to Ismail-type integrals: for $n\in\mathbb{N}_{0}$,
$\displaystyle\frac{1}{2}\displaystyle\int_{-\infty}^{\infty}\displaystyle\frac{x^{n+1}dx}{\cos(\sqrt{x}K)-\cosh(\sqrt{x}K^{\prime})}=(-1)^{n+1}4\displaystyle\frac{d^{2n+1}}{du^{2n+1}}\displaystyle\frac{{\rm sn}^{2}(u,k)}{{\rm cd}^{2}(u,k){\rm sd}(2u,k)}|_{u=0},$ (1.5)
@@
Furthermore, we put forward the specific coefficients preceding each term and set up some explicit relations between the Berndt-type integrals and Barnes multiple zeta functions. Additionally, we establish their relationship with Barnes multiple zeta functions, offering new closed-form evaluations and insights (see Theorem 7.2).
@@ 2 Some Preliminary Results
$x:=k^{2},\ y\left(x\right):=\pi K^{\prime}/K,\ q\equiv q\left(x\right):=e^{-y},\ z:=z\left(x\right)=2K/\pi,\ z^{\prime}=dz/dx,$ (2.9)
@@
$\,{}_{2}F_{1}\left(a,b,\displaystyle\frac{a+b+1}{2};\displaystyle\frac{1}{2}\right)=\displaystyle\frac{\Gamma\left(\displaystyle\frac{1}{2}\right)\Gamma\left(\displaystyle\frac{a+b+1}{2}\right)}{\Gamma\left(\displaystyle\frac{a+1}{2}\right)\Gamma\left(\displaystyle\frac{b+1}{2}\right)},$
@@
$\displaystyle\frac{d^{n}z}{dx^{n}}\mid_{x=\tfrac{1}{2}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}\sqrt{\pi}}{\Gamma^{2}\left(\displaystyle\frac{n}{2}+\displaystyle\frac{3}{4}\right)}.$ (2.10)
In particular, taking $x=\displaystyle\frac{1}{2}$ and $n=0,1,2$ in (2.10), we obtain that
@@
which are essential for our explicit evaluations. The derivative relation:
$\frac{dx}{dy}=-x(1-x)z^{2}$ (2.11)
@@ 3 Some Definitions and Lemmas
Let $s\in\mathbb{C}$ and $a,b,\theta\in\mathbb{R}$ with $|\theta|<2b\pi$ and $a,b\neq 0$. Define
$Z_{1}(s,\theta;a,b)$ $:=\displaystyle\frac{\pi^{4}\sinh(\theta s)}{\sin(a\pi s)\sinh(b\pi s)\cosh^{2}(b\pi s)},$ $Z_{2}(s,\theta;a,b):$ $=\displaystyle\frac{\pi^{4}\cosh(\theta s)}{\cos(a\pi s)\sinh(b\pi s)\cosh^{2}(b\pi s)},$
$Z_{3}(s,\theta;a,b)$ $:=\displaystyle\frac{\pi^{4}\sinh(\theta s)}{\cos(a\pi s)\sinh^{2}(b\pi s)\cosh(b\pi s)},$ $Z_{4}(s,\theta;a,b):$ $=\displaystyle\frac{\pi^{4}\sinh(\theta s)}{\cos(a\pi s)\cosh^{3}(b\pi s)}.$
@@
Let $m\in\mathbb{N}$ and $p\in\mathbb{Z}$. Define
$G_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh(ny)\cosh^{m}(ny)},\quad X_{p,m}(y):=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh^{m}(ny)},$
$G_{p,m}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}}{\sinh((2n-1)y/2)\cosh^{m}((2n-1)y/2)},$
$\overline{G}_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\sinh^{m}((2n-1)y/2)\cosh((2n-1)y/2)},$
$Y_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p-1}}{\sinh^{m}((2n-1)y/2)},$
$\mathrm{DY}_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}\cosh((2n-1)y/2)}{\sinh^{m}((2n-1)y/2)},$
$X_{p,m}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\cosh^{m}((2n-1)y/2)},$
$\mathrm{DX}_{p,m}^{\prime}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p+1}\sinh((2n-1)y/2)}{\cosh^{m}((2n-1)y/2)},$
$B_{p,m}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p-1}}{\cosh^{m}(ny)},\quad\mathrm{DB}_{p,m}(y):=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}\sinh(ny)}{\cosh^{m}{ny}}.$
@@
(Residue Theorem [9]) Let $\xi(s)$ be a kernel function and let $r(s)$ be a function which is $O(s^{-2})$ at infinity. Then
$\sum_{\begin{subarray}{c}\alpha\in O\\ \end{subarray}}\text{Res}\left(r(s)\xi(s),s=\alpha\right)+\displaystyle\sum_{\begin{subarray}{c}\beta\in S\\ \end{subarray}}\text{Res}\left(r(s)\xi(s),s=\beta\right)=0,$ (3.12)
where $S$ is the set of poles of $r(s)$ and $O$ is the set of poles of $\xi(s)$ that are not poles $r(s)$. Here $\text{Res}\left(r(s),s=\alpha\right)$ denotes the residue of $r(s)$ at $s=\alpha$. The kernel function $\xi(s)$ is meromorphic in the whole complex plane and satisfies $\xi(s)=o(s)$ over an infinite collection of circles $|s|=\rho_{k}$ with $\rho_{k}\rightarrow\infty$.
@@
([23, Thm. 2.4]) Let $x,y,z$ and $z^{\prime}$ satisfy (2.9). Given the formula $\Omega(x,e^{-y},z,z^{\prime})=0$, we have the transformation formula
@@
(cf.[20, Lem. 1.2]) Let $n$ be an integer, then the following formulas hold:
$\displaystyle\frac{\pi}{\sin(\pi s)})\overset{s\rightarrow n}{=}(-1)^{n}\left(\displaystyle\frac{1}{s-n}+2\displaystyle\sum_{k=1}^{\infty}\overline{\zeta}(2k)(s-n)^{2k-1}\right),$ (3.14)
$\displaystyle\frac{\pi}{\sinh(\pi s)}\overset{s\rightarrow ni}{=}(-1)^{n}\left(\displaystyle\frac{1}{s-ni}+2\displaystyle\sum_{k=1}^{\infty}(-1)^{k}\overline{\zeta}(2k)(s-ni)^{2k-1}\right),$ (3.15)
$\displaystyle\frac{\pi}{\cos(\pi s)}\overset{s\rightarrow n-1/2}{=}(-1)^{n}\left\{\displaystyle\frac{1}{s-\displaystyle\frac{2n-1}{2}}+2\displaystyle\sum_{k=1}^{\infty}\overline{\zeta}(2k)\left(s-\displaystyle\frac{2n-1}{2}\right)^{2k-1}\right\},$ (3.16)
$\displaystyle\frac{\pi}{\cosh(\pi s)}\overset{s\rightarrow(n-1/2)i}{=}(-1)^{n}i\left\{\displaystyle\frac{1}{s-\displaystyle\frac{2n-1}{2}i}+2\displaystyle\sum_{k=1}^{\infty}(-1)^{k}\overline{\zeta}(2k)\left(s-\displaystyle\frac{2n-1}{2}i\right)^{2k-1}\right\}.$ (3.17)
Lemma 3.4.
([21, Eqs. (4.4)-(4.7)]) Let $n\in\mathbb{Z}$. We have
$\displaystyle\frac{\left(-1\right)^{n}}{\sin\left(\displaystyle\frac{1\pm i}{2}z\right)}\overset{z\rightarrow\left(1\mp i\right)n\pi}{=}\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)n\pi}+2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)n\pi\right)^{2k-1}},$ (3.18)
$\displaystyle\frac{\left(-1\right)^{n-1}}{\cos\left(\displaystyle\frac{1\pm i}{2}z\right)}\overset{z\rightarrow\left(1\mp i\right)\tilde{n}\pi}{=}\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)\tilde{n}\pi}-2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)\tilde{n}\pi\right)^{2k-1}},$ (3.19)
where $\tilde{n}:=n-\displaystyle\frac{1}{2}$. Here $\zeta(s)$ and $\overline{\zeta}(s)$ stand for the Riemann zeta function and alternating Riemann zeta function, respectively, which are defined by
@@
When $s=2m$ is an even, Euler proved the famous formula
@@
where $\mathrm{nd}(u):=1/\mathrm{dn}(u)$, $\mathrm{S}_{2n}\left(x\right),\mathrm{A}_{2n}\left(x\right)\in\mathbb{Z}[x].$

### v3

@@
Abstract. In this paper, we focus on the calculation of a specific type of Berndt integral, which exclusively involves (hyperbolic) cosine functions. Initially, this integral is transformed into a Ramanujan-type hyperbolic (infinite) sum via contour integration. Subsequently, a function incorporating $\theta$ is defined. By employing the residue theorem, the mixed Ramanujan-type hyperbolic (infinite) sum with both hyperbolic cosine and hyperbolic sine in the denominator is converted into a simpler Ramanujan-type hyperbolic (infinite) sum, which contains only hyperbolic cosine or hyperbolic sine in the denominator. The simpler Ramanujan-type hyperbolic (infinite) sum is then evaluated using Jacobi elliptic functions, Fourier series expansions, and Maclaurin series expansions. Ultimately, the result is expressed as a rational polynomial of $\Gamma(1/4)$ and $\pi^{-1/2}$. Additionally, the integral is related to the Barnes multiple zeta function, which provides an alternative method for its calculation.
@@ 1 Introduction
where $s\geq 1$ and $m\geq 1$ for the “++” case and $s\geq 2m+1$ otherwise. In fact, the study of Berndt-type integrals can be traced back to 1912, when the renowned Indian mathematician Ramanujan investigated the following improper integrals involving trigonometric and hyperbolic functions. He conjectured the integral identity (see [16, pp. 325-326])
@@
Here $K\equiv K(x)$ denotes the complete elliptic integral of the first kind and $K^{\prime}:=K(1-x)$. Kuznetsov [14] generalized Ismail-type integrals using contour integration and theta functions, proving
$\displaystyle\frac{1}{2}\displaystyle\int_{-\infty}^{\infty}\displaystyle\frac{t^{n}\mathrm{d}t}{\cos(K\sqrt{t})+\cosh(K^{\prime}\sqrt{t})}=(-1)^{n}\displaystyle\frac{\mathrm{d}^{2n+1}}{\mathrm{d}u^{2n+1}}\displaystyle\frac{{\rm sn}(u,k)}{{\rm cd}(u,k)}|_{u=0},$ (1.3)
@@
In the past decade, research on these analogous integrals has attracted increasing interest, yielding notable progress. Recent advances in this direction can be found in the works of [9, 15, 18, 19, 23, 24, 26, 27] and the references therein. A significant turning point in this line of research occurred in 2016 when Professor Berndt [6] developed a systematic approach to study ${\rm BI}^{\pm}(s,1)$. Berndt further advanced the field by evaluating such integrals directly using Cauchy’s residue theory, Fourier series expansions, and Maclaurin series expansions of Jacobi elliptic functions. In their series of works [19, 23, 24], Rui, Xu, and Zhao developed an extended framework based on Berndt’s approach, which led to fundamental structural theorems for higher-order Berndt-type integrals. Letting $X=\Gamma^{4}(1/4)$ and $Y=\pi^{-1}$, they proved that for all integers $m\geq 1$ and $p\geq[m/2]$ (see [24, Thms 1.1 and 1.2]),
@@
Very recently, a groundbreaking connection between Berndt-type integrals and Barnes multiple zeta functions (See Section 7) was established by Bradshaw and Vignat [9, Cor. 5], who also linked Jacobi elliptic functions to Ismail-type integrals: for $n\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$,
$\displaystyle\frac{1}{2}\displaystyle\int_{-\infty}^{\infty}\displaystyle\frac{x^{n+1}\mathrm{d}x}{\cos(\sqrt{x}K)-\cosh(\sqrt{x}K^{\prime})}=(-1)^{n+1}4\displaystyle\frac{\mathrm{d}^{2n+1}}{\mathrm{d}u^{2n+1}}\displaystyle\frac{{\rm sn}^{2}(u,k)}{{\rm cd}^{2}(u,k){\rm sd}(2u,k)}|_{u=0},$ (1.5)
@@
Furthermore, we compute the specific coefficients preceding each term and set up some explicit relations between the Berndt-type integrals and Barnes multiple zeta functions. Additionally, we establish their relationship with Barnes multiple zeta functions, offering new closed-form evaluations and insights (see Theorem 7.3). Furthermore, it should be noted that the last two authors of this paper, together with Chen in a recent joint work [7], applied methods similar to those used in the present article to obtain a structural result analogous to (1) for a mixed Berndt-type integral. Specifically, in [1], the denominator in (1.7) was replaced by $\bigl[\cosh(2x)-\cos(2x)\bigr]\bigl[\cosh x-\cos x\bigr],$ and the numerator exponent was taken as s=4m3s=4m-3.
@@ 2 Some Preliminary Results
$x:=k^{2},\ y\left(x\right):=\pi K^{\prime}/K,\ q\equiv q\left(x\right):=e^{-y},\ z:=z\left(x\right)=2K/\pi,\ z^{\prime}=\mathrm{d}z/\mathrm{d}x,$ (2.9)
@@
$_{2}F_{1}\left(a,b,\displaystyle\frac{a+b+1}{2};\displaystyle\frac{1}{2}\right)=\displaystyle\frac{\Gamma\left({1}/{2}\right)\Gamma\left({(a+b+1)}/{2}\right)}{\Gamma\left({(a+1)}/{2}\right)\Gamma\left({(b+1)}/{2}\right)},$
@@
$\displaystyle\frac{\mathrm{d}^{n}z}{\mathrm{d}x^{n}}\mid_{x=\tfrac{1}{2}}=\displaystyle\frac{\left(1/2\right)_{n}^{2}\sqrt{\pi}}{\Gamma^{2}\left(\displaystyle\frac{n}{2}+\displaystyle\frac{3}{4}\right)}.$ (2.10)
In particular, taking $n=0,1,2$ in (2.10), we obtain that
@@
which are essential for our explicit evaluations. The derivative relation (see [5, pp. 120, Entry 9(i)]):
$\frac{\mathrm{d}x}{\mathrm{d}y}=-x(1-x)z^{2}$ (2.11)
@@ 3 Some Definitions and Lemmas
Let $s\in\mathbb{C}$, $g_{1},g_{2},g_{3},g_{4}\in\mathbb{N}_{0}$ and $a,b,\theta\in\mathbb{R}$ with $|\theta|<2b\pi$ and $a,b\neq 0$. Define
$Z(s,\theta;a,b|g_{1},g_{2},g_{3},g_{4}):=\displaystyle\frac{\pi^{g_{1}+g_{2}+g_{3}+g_{4}}\sinh(\theta s)}{\sin^{g_{1}}(a\pi s)\cos^{g_{2}}(a\pi s)\sinh^{g_{3}}(b\pi s)\cosh^{g_{4}}(b\pi s)},$
$\bar{Z}(s,\theta;a,b|g_{1},g_{2},g_{3},g_{4}):=\displaystyle\frac{\pi^{g_{1}+g_{2}+g_{3}+g_{4}}\cosh(\theta s)}{\sin^{g_{1}}(a\pi s)\cos^{g_{2}}(a\pi s)\sinh^{g_{3}}(b\pi s)\cosh^{g_{4}}(b\pi s)}.$
@@
Let $m_{1},m_{2}\in\mathbb{Z}$, $m_{1}+m_{2}\geq 1$ and $p\in\mathbb{Z}$. Define
$G_{p,m_{1},m_{2}}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}n^{p}}{\sinh^{m_{1}}(ny)\cosh^{m_{2}}(ny)},$
$\tilde{G}_{p,m_{1},m_{2}}(y)$ $:=\displaystyle\sum_{n=1}^{\infty}\displaystyle\frac{(-1)^{n}(2n-1)^{p}}{\sinh^{m_{1}}((2n-1)y/2)\cosh^{m_{2}}((2n-1)y/2)}.$
@@
([11]) Let $\xi(s)$ be a kernel function and let $r(s)$ be a function that is $\mathscr{O}(s^{-2})$ at infinity. Then
$\sum_{\begin{subarray}{c}\alpha\in O\\ \end{subarray}}\mathrm{Res}\left(r(s)\xi(s),s=\alpha\right)+\displaystyle\sum_{\begin{subarray}{c}\beta\in S\\ \end{subarray}}\mathrm{Res}\left(r(s)\xi(s),s=\beta\right)=0,$ (3.12)
where $S$ is the set of poles of $r(s)$ and $O$ is the set of poles of $\xi(s)$ that are not poles of $r(s)$. Here $\text{Res}\left(r(s),s=\alpha\right)$ denotes the residue of $r(s)$ at $s=\alpha$. The kernel function $\xi(s)$ is meromorphic in the whole complex plane and satisfies $\xi(s)=o(s)$ over an infinite collection of circles $|s|=\rho_{k}$ with $\rho_{k}\rightarrow\infty$.
@@
([25, Thm. 2.4]) Let $x,y,z$ and $z^{\prime}$ satisfy (2.9). Given the formula $\Omega(x,e^{-y},z,z^{\prime})=0$, we have the transformation formula
@@
([22, Lem. 1.2]) Let $n$ be an integer, then the following Laurent expansions hold:
$\displaystyle\frac{\pi}{\sin(\pi s)}=(-1)^{n}\left(\displaystyle\frac{1}{s-n}+2\displaystyle\sum_{k=1}^{\infty}\overline{\zeta}(2k)(s-n)^{2k-1}\right),$ (3.14)
$\displaystyle\frac{\pi}{\sinh(\pi s)}=(-1)^{n}\left(\displaystyle\frac{1}{s-ni}+2\displaystyle\sum_{k=1}^{\infty}(-1)^{k}\overline{\zeta}(2k)(s-ni)^{2k-1}\right),$ (3.15)
$\displaystyle\frac{\pi}{\cos(\pi s)}=(-1)^{n}\left\{\displaystyle\frac{1}{s-\displaystyle\frac{2n-1}{2}}+2\displaystyle\sum_{k=1}^{\infty}\overline{\zeta}(2k)\left(s-\displaystyle\frac{2n-1}{2}\right)^{2k-1}\right\},$ (3.16)
$\displaystyle\frac{\pi}{\cosh(\pi s)}=(-1)^{n}i\left\{\displaystyle\frac{1}{s-\displaystyle\frac{2n-1}{2}i}+2\displaystyle\sum_{k=1}^{\infty}(-1)^{k}\overline{\zeta}(2k)\left(s-\displaystyle\frac{2n-1}{2}i\right)^{2k-1}\right\}.$ (3.17)
Here $\zeta(s)$ and $\overline{\zeta}(s)$ stand for the Riemann zeta function and alternating Riemann zeta function, respectively, which are defined by
@@
When $s=2m$ is even, Euler proved the famous formula
@@
All these Laurent expansions have non-empty domain (annulus) of convergence. As these functions are meromorphic (with only poles as singularities), they are analytic in the complex plane except at the poles. Therefore, there are no singularities in the annular region r<|sa|<Rr<|s-a|<R(for any expansion center $a$, $r$ as the distance from the center $a$ to its nearest singularity and $R$ as the distance from the center $a$ to its second nearest singularity), which satisfies the convergence condition of the Laurent series. What’s more, for any integer $n$, the distance $r$ from the center $a$ to its nearest singularity is positive (since the poles are discrete and have no accumulation points), and $r<R$ (the distance from the center to the second nearest singularity is greater than that to the nearest one). Thus, the annular region r<|sa|<Rr<|s-a|<R is non-empty.
Lemma 3.4.
([23, Eqs. (4.4)-(4.7)]) Let $n\in\mathbb{Z}$. We have
$\displaystyle\frac{\left(-1\right)^{n}}{\sin\left(\displaystyle\frac{1\pm i}{2}z\right)}=\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)n\pi}+2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)n\pi\right)^{2k-1}},$ (3.18)
$\displaystyle\frac{\left(-1\right)^{n-1}}{\cos\left(\displaystyle\frac{1\pm i}{2}z\right)}=\displaystyle\frac{1\mp i}{z-\left(1\mp i\right)\tilde{n}\pi}-2\displaystyle\sum_{k=1}^{\infty}{\displaystyle\frac{\overline{\zeta}\left(2k\right)}{\pi^{2k}}\left(\displaystyle\frac{1\pm i}{2}\right)^{2k-1}\left(z-\left(1\mp i\right)\tilde{n}\pi\right)^{2k-1}},$ (3.19)
where $\tilde{n}:=n-\displaystyle\frac{1}{2}$. As in Lemma 3.3, since the singularities of a meromorphic function are discrete, we can always find an annular region centered at the expansion center, which lies between the nearest and second nearest singularities. This ensures that the Laurent expansion converges in this region and guarantees a non-empty convergence domain.
@@
where $\mathrm{nd}(u):=1/\mathrm{dn}(u)$, and $\mathrm{S}_{2n}\left(x\right),\mathrm{A}_{2n}\left(x\right)\in\mathbb{Z}[x]$. Here $x=k^{2}$ is consistent with the notation used in (2.9).